OpenStudy (anonymous):

can someone help me with some math questions such as this square root 18x* square root 6x^3 helpppppppppppppp please

OpenStudy (jdoe0001):

\(\bf \sqrt{18x}\cdot \sqrt{6x^3}\implies \sqrt{18x^1}\cdot \sqrt{6x^3}\implies \sqrt{18x^1\cdot 6x^3}\) keep in mind that \(\bf \large a^{\color{blue}{ n}}\cdot a^{\color{blue}{ m}}\implies a^{{\color{blue}{ n+m}}}\)

OpenStudy (jdoe0001):

if the factors have the same base, different exponent keep the base, sum the exponents

OpenStudy (anonymous):

so what would the answer be jdoe0001